OpenStudy (anonymous):

If the margin of error in an estimate for the mean weight of a shipment is + or -2 pounds at a confidence level of 95 percent, what will be the margin of error at a confidence level of 98 percent? Be sure to show how you arrived at your answer.

OpenStudy (perl):

we can use that equation we used earlier

OpenStudy (perl):

x bar + - margin of error. margin of error = z* st.dev / sqrt(n)

OpenStudy (anonymous):

i got it haha

OpenStudy (perl):

z* for 95% confidence level is 1.960 z* for 98% confidence is 2.326 so we are given that 1.960 * st.dev / sqrt(n) = 2

OpenStudy (anonymous):

Thanks! @perl

OpenStudy (perl):

what did you get for the answer

OpenStudy (anonymous):

i got 2

OpenStudy (perl):

st.dev / sqrt(n) = 2/ 1.960 so 2.326 * 2/ 1.960 = 2.37 ,

OpenStudy (anonymous):

i rounded haha but i got something close i got 2.47

OpenStudy (perl):

should be 2.37 pounds

OpenStudy (perl):

should we round?

OpenStudy (anonymous):

yes no haha

OpenStudy (perl):

yes no?

OpenStudy (perl):

yes no sounds like a maybe ;)

OpenStudy (anonymous):

You're the best @perl

OpenStudy (perl):

well i dont think the answer is 2 because then the confidence level didn't change, right?

OpenStudy (anonymous):

maybe you can help me with some more stats if ever do get stumped?

OpenStudy (perl):

sure :)

OpenStudy (perl):

i would love that